# Addressing Sample Inefficiency in Multi-View Representation Learning

### v1

@@ C.3 Explaining training dynamics in low patch sampling regime
We now turn to a simple form of the augmentation graph to understand how using low number of augmentations affects the evolution of $ZZ^{T}$. Minimizing eq. 29 implies that the spectral decomposition of $Z$ would align with the top eigenvectors (and values) of $\overline{\mathcal{W}}$. We will demonstrate that in the low sampling regime (using few augmentations), the eigenvectors of the sampled augmentation graph $\tilde{\mathcal{W}}$ may not align with those of $\overline{\mathcal{W}}$.
Augmentation graph setup. We define an augmentation graph with only two instances from two different classes, similar to the one presented in [27]. Let us denote the four instances as $\bar{x}_{i}$ for $i\in{1,2,3,4}$, where $\bar{x}_{1},\bar{x}_{2}$ belong to class 1 (i.e. $y_{1},y_{2}=1$) and $\bar{x}_{3},\bar{x}_{4}$ belong to class 2 (i.e. $y_{3},y_{4}=4$). Let us further assume that $\bar{x}_{1},\bar{x}_{3}$ have the highest pixel-level similarity among $(\bar{x}_{1},\bar{x}_{i})\forall i\in{2,3,4}$, thereby making it more likely to have similar patches (see Figure 2 for illustration). We denote this relationship among input examples using $\mathcal{G}$ to indicate (pixel-wise) global similarity groups. So, $\mathcal{G}_{1},\mathcal{G}_{3}=1$ and $\mathcal{G}_{2},\mathcal{G}_{4}=2$. We can use the following probabilistic formulation to model our augmentation functions (see Figure 2B):
@@ Appendix D Implementation Details
Image Classification Datasets Across all experiments, our settings mainly follow [6]. In particular, we run Table 1 summarizes our pretraining settings on Cifar-10 [21], and STL-10 [9]. In Table 1, we outline the corresponding linear evaluation settings for Resnet-50. Note that we add a linear classifier layer to the encoder’s features and discard the projection layers for evaluation.
(a) Pretraining
@@
batch size 128
epochs 100 (CIFAR10), 50 (STL10)
@@
D.1 Longer Pretraining
Algorithm Best accuracy Best accuracy @ epoch
Barlow-Twins (2-augs) w/ pdim=256 92.04 +/- 0.16 400
Barlow-Twins (4-augs) w/ pdim=256 92.39 +/- 0.17 340
Barlow-Twins (8-augs) w/ pdim=256 92.64 +/- 0.10 140
D.2 SwAV-like augmentations for compute efficient multi-augmentation framework
D.3 On training with full dataset with 4/8 augmentations
Figure 10: BarlowTwins pretraining on full CIFAR-10 dataset.
Algorithm #augs=2 #augs=4 #augs=8
Barlow-Twins w/ pdim=256 86.43 +/- 0.72 91.73 +/- 0.16 92.71 +/- 0.19
Barlow-Twins w/ pdim=8192 85.44 +/- 0.54 91.40 +/- 0.32 92.40 +/- 0.13
D.4 Empirical results on transfer learning

### v2

@@ C.3 Explaining training dynamics in low patch sampling regime
We now turn to a simple form of the augmentation graph to understand how using low number of augmentations affects the evolution of $ZZ^{T}$. Minimizing Equation 41 implies that the spectral decomposition of $Z$ would align with the top eigenvectors (and values) of $\overline{\mathcal{W}}$. We will demonstrate that in the low sampling regime (using few augmentations), the eigenvectors of the sampled augmentation graph $\tilde{\mathcal{W}}$ may not align with those of $\overline{\mathcal{W}}$.
Augmentation graph setup. We define an augmentation graph with only two instances from two different classes, similar to the one presented in [35]. Let us denote the four instances as $\bar{x}_{i}$ for $i\in{1,2,3,4}$, where $\bar{x}_{1},\bar{x}_{2}$ belong to class 1 (i.e. $y_{1},y_{2}=1$) and $\bar{x}_{3},\bar{x}_{4}$ belong to class 2 (i.e. $y_{3},y_{4}=4$). Let us further assume that $\bar{x}_{1},\bar{x}_{3}$ have the highest pixel-level similarity among $(\bar{x}_{1},\bar{x}_{i})\forall i\in{2,3,4}$, thereby making it more likely to have similar patches. We denote this relationship among input examples using $\mathcal{G}$ to indicate (pixel-wise) global similarity groups. So, $\mathcal{G}_{1},\mathcal{G}_{3}=1$ and $\mathcal{G}_{2},\mathcal{G}_{4}=2$. We can use the following probabilistic formulation to model our augmentation functions (see Figure 6B):
@@ Appendix D Implementation Details
Image Classification Datasets Across all experiments, our settings mainly follow [13]. In particular, Table 4 summarizes our pretraining settings on Cifar-10 [26], STL-10 [14] and Imagenet-100 [33]. The Imagenet-100 dataset was generated by sampling 100 classes from the original Imagenet-1k dataset, according to this list [38]. In Table 4, we outline the corresponding linear evaluation settings for Resnet-50 (for CIFAR-10 and STL-10) and ResNet-18 (for Imagenet). Note that we add a linear classifier layer to the encoder’s features and discard the projection layers for evaluation. Our code base is publicly available on github.
@@
batch size 128 (Imagnet), 256 (CIFAR, STL)
epochs 100
@@
The key SSL loss functions that we use in this work are BarlowTwins [43] and VICReg [5]. Let us suppose that the embeddings of two augmentations of a batch of images are denoted as $z$ and $z^{\prime}$. The BarlowTwins loss function is as follows:
$\mathcal{L}_{BT}$ $=\sum_{i}(C_{ii}-1)^{2}+\beta\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (47)
where $\text{where }\quad C$ $=\frac{1}{n-1}\sum_{k=1}^{n}(z_{k}-\bar{z})(z_{k}^{\prime}-\bar{z^{\prime}})^{T}$
and $\text{and }\quad\bar{z}$ $=\frac{1}{n}\sum_{k=1}^{n}z_{k}\quad\text{,}\quad\bar{z^{\prime}}=\frac{1}{n}\sum_{k=1}^{n}z_{k}^{\prime}$ (48)
$C_{ij}$ is the element of $C$ at row $i$, column $j$ and $n$ is the batch size. For each projector dimensionality, $d$, we search for the hyperparameter, $\beta$, that yields the best downstream task performance.
The VICReg loss function is as follows:
$\mathcal{L}_{VIC}$ $=\frac{1}{n}\mu\sum_{k=1}^{n}\|z_{k}-z_{k}^{\prime}\|^{2}+\frac{1}{2}\mu\left[v(Z)+v(Z^{\prime})\right]+\frac{1}{2}\left[c(Z)+c(Z^{\prime})\right]$ (49)
where $\text{where }\quad v(Z)$ $=\frac{1}{d}\sum_{i=1}^{d}max(0,1-Stdev(z_{:,i}))$
and $\text{and }\quad c(Z)$ $=\frac{1}{d}\sum_{i}\sum_{j\neq i}\left[C(Z)_{ij}\right]^{2}\quad\text{,}\quad C(Z)=\frac{1}{n-1}\sum_{k=1}^{n}(z_{k}-\bar{z})(z_{k}-\bar{z})^{T}$
For each projector dimensionality, $d$, we search for the hyperparameter, $\mu$, that yields the best downstream task performance.
D.1 Empirical results for low-dimensional projectors
Figure 8: Low-dimensional projectors can yield good representations for both BarlowTwins and VICReg. We demonstrate that using a higher orthogonality constraint, $\beta$, for lower projector dimensionality can achieve similar performance over a wide range of projector dimensions ($d$). Note that for VICReg, we plot the ratio of the coefficient of the covariance loss to the coefficient of the invariance loss, i.e. $\beta=\frac{1}{d*\mu}$, where $\mu$ is the coefficient of the invariance loss. (See Equation 49 for details of the loss formulation.)
pdim Projector params (approx) Barlow Twins VICReg
fixed $\beta$ optimal $\beta^{*}$ fixed $\beta$ optimal $\beta^{*}$
64 135k 73.6 $\pm$ 0.9 82.1 $\pm$ 0.2 68.9 $\pm$ 0.2 81.9 $\pm$ 0.1
128 278k 74.7 $\pm$ 1.4 83.0 $\pm$ 1.1 70.6 $\pm$ 0.3 82.3 $\pm$ 0.4
256 589k 75.9 $\pm$ 0.7 83.4 $\pm$ 0.4 75.3 $\pm$ 0.2 81.9 $\pm$ 0.3
512 1.3M 79.2 $\pm$ 0.8 82.8 $\pm$ 0.5 79.3 $\pm$ 0.4 82.1 $\pm$ 0.6
1024 3.1M 81.3 $\pm$ 1.0 82.9 $\pm$ 0.3 79.2 $\pm$ 0.9 82.5 $\pm$ 0.9
2048 8.3M 81.0 $\pm$ 0.9 82.3 $\pm$ 0.5 80.6 $\pm$ 0.0 81.9 $\pm$ 1.2
4096 25.2M 82.3 $\pm$ 0.4 82.3 $\pm$ 0.4 80.5 $\pm$ 0.3 81.0 $\pm$ 0.4
8192 83.9M 82.2 $\pm$ 0.4 82.2 $\pm$ 0.4 80.4 $\pm$ 1.5 80.4 $\pm$ 1.5
Table 5: Extended version of Table 1. Optimizing for orthogonality appropriately allows low-dimensional projectors to match the performance for BarlowTwins and VICReg (on CIFAR-10) of much higher-dimensional projectors.
D.2 Empirical results with multi-augmentations along with Time
Figure 9: Using multiple augmentations improves representation learning performance and convergence. (A-C) Across BarlowTwins and VICReg for CIFAR-10 and STL-10 pretraining, using 4 augmentations instead of 2 helps improve performance. (D-F) Although the 4-augmentations take longer for each epoch, its performance still trumps the 2-augmentation version of the algorithm at the same wall clock time. Please see Section E.3 for more results.
Figure 10: Multi-augmentation improves sample efficiency, recovering similar performance with significantly fewer unique samples in the pretraining dataset. Across BarlowTwins and VICReg pretraining on CIFAR-10 and STL-10, for the same effective dataset size (#$\#augs\times\#unique\_samples$), using more patches improves performance at the same epoch (A-C) or wall clock time (D-F). However, a tradeoff exists wherein more data augmentations fail to improve performance in the scarce data regime.
Figure 11: BarlowTwins pretraining on full Imagenet-100 dataset with 2, 4 and 8 augmentations.
Figure 12: BarlowTwins pretraining on fraction of Imagenet-100 dataset with 2, 4 and 8 augmentations.
D.3 Empirical results on transfer learning
@@
Appendix E Additional Experiments probing multi-augmentation learning
E.1 Longer Pretraining to determine early stopping
Algorithm Best accuracy Best accuracy @ epoch
Barlow-Twins (2-augs) w/ pdim=256 92.04 +/- 0.16 400
Barlow-Twins (4-augs) w/ pdim=256 92.39 +/- 0.17 340
Barlow-Twins (8-augs) w/ pdim=256 92.64 +/- 0.10 140
E.2 SwAV-like augmentations for compute efficient multi-augmentation framework
E.3 Training with full dataset with 4/8 augmentations
Figure 21: BarlowTwins pretraining on full CIFAR-10 dataset with 2, 4 and 8 augmentations.
Algorithm #augs=2 #augs=4 #augs=8
Barlow-Twins w/ pdim=256 86.43 +/- 0.72 91.73 +/- 0.16 92.71 +/- 0.19
Barlow-Twins w/ pdim=8192 85.44 +/- 0.54 91.40 +/- 0.32 92.40 +/- 0.13